OpenStudy (anonymous):

Let f(x) = ln(8 − x).Find f−1(x).

OpenStudy (tyteen4a03):

Hint: \(e^x = a\) = \(ln(a) = x\)

OpenStudy (anonymous):

e^x=e^(2y + In(y))

OpenStudy (tyteen4a03):

@gmurillo What is this?

OpenStudy (anonymous):

I was looking at the wrong notes. Is the answer e^y -8=x

OpenStudy (tyteen4a03):

Correct. :)

OpenStudy (anonymous):

thanks. Do you know how to find the range?

OpenStudy (tyteen4a03):

The range of which equation?

OpenStudy (anonymous):

of the original equation

OpenStudy (amoodarya):

if you mean inverse of that function look the draw